OpenStudy (lgbasallote):

is \[\huge \frac{- (-1)^n}{n! (s+a)^2} = \frac{1}{(s+a)^{n+1}}\] if so..how can i show it?

OpenStudy (lgbasallote):

@sami-21 please abandon my very difficult question and help me here instead :DDD hehe

OpenStudy (anonymous):

try it with some arbitrary values

OpenStudy (klimenkov):

Put \(n=1\), \(s=1\) and \(a=0\) and check identity.

OpenStudy (lgbasallote):

maybe i should state my solution and how i got this?

OpenStudy (lgbasallote):

basically..i was trying to prove \[\mathcal L \{ \frac{t^n e^{-at}}{n!} \} = \frac{1}{(s+a)^{n+1}}\]using the formula \[\mathcal L \{t^n f(t) \} = (-1)^n \frac{d^n}{ds^n} (\mathcal L \{f(t) \} )\]

OpenStudy (lgbasallote):

cute symbols got into cute question now :D

OpenStudy (anonymous):

ahh laplace transforms

OpenStudy (lgbasallote):

yeahh

OpenStudy (lgbasallote):

maybe i did a wrong step? or maybe i shouldnt have used that formula?

OpenStudy (anonymous):

do you know what laplace is?

OpenStudy (anonymous):

laplace transform is \[\int f(t)e^{-st}dt=F(s)\]

OpenStudy (lgbasallote):

although.. i do know \[\mathcal L \{e^{-at} t^n \} = \frac{n!}{(s+a)^{n+1}}\] but i still want to prove it some other way...

OpenStudy (lgbasallote):

and yes im familiar of teh general form (if that's what it's called)

OpenStudy (anonymous):

so to prove i think you'd have to do \[\int \frac{e^{-st}t^ne^{-at}}{n!}\]dt

OpenStudy (anonymous):

\[\int \frac{e^{(-s-a)t}t^n}{n!}dt\]

OpenStudy (lgbasallote):

so i cant use \[\large \mathcal L \{ t f(t) \} = (-1)^n \frac{d^n}{ds^n} (F(s) )?\] i can only use the "general equation"?

OpenStudy (lgbasallote):

that should be L{t^n f(t)}

OpenStudy (anonymous):

wouldn't that be usng another proof to prove another proof? proofception?

OpenStudy (lgbasallote):

well im not really asked to prove it...im practicing my skills by proving the table in my book

OpenStudy (anonymous):

best way would to use the definition of laplace's transform. however the reason why the book just gives you those is because the proof my be a huge problem... i'd check my book for the proof but it's outside=/

OpenStudy (anonymous):

all of these transforms on the table derive from the laplace transform definition

OpenStudy (lgbasallote):

well im just practicing my use and mastery of that formula i wrote...so i was asking if it was applicable

OpenStudy (anonymous):

it's possible

OpenStudy (lgbasallote):

so is my result right? how can i change it to the other form?

OpenStudy (lgbasallote):

is it too hard? should i move on and come back to it later on?

OpenStudy (anonymous):

so let me see you divided by n! and it was put into the laplace... i don't think that's right

OpenStudy (anonymous):

i'm not 100% sure but ithink it should be the laplace of the top over n!

OpenStudy (anonymous):

which would make sense because

OpenStudy (anonymous):

if you have n! as a denominator it'd cancel with the top to get one

OpenStudy (lgbasallote):

yes exactly

OpenStudy (anonymous):